### v1

@@ VI Discussion
This study proposed a novel framework that can ensure the safety of automation for robotic surgery in a theoretically rigorous way. Our results were demonstrated in a scenario based on ESD, a representative endoscopic surgical method for the treatment of early GI cancer, by using an endoscopic robot. The proposed framework utilizes CBFs to enable the automated motion of the robot, allowing it to clearly define the boundaries of each tumor, even when multiple tumors are in close proximity within the GI tract. This allows individual removal of each tumor while enforcing safety guarantees on the surrounding normal mucosal tissues.
@@
A model-free control scheme was leveraged, thus avoiding the complexity of robot dynamics modeling. However, even though a model-free control scheme was used, the simulation environment can differ from the actual environment where the robotic system is implemented. From (24), the proposed framework employs a simple model-free velocity controller to track the safe velocity $\dot{\bm{x}}_{s}$. Here, since $\dot{\bm{e}}$ is expressed as $J^{+}(\bm{q})(\dot{\bm{x}}-\dot{\bm{x}}_{s})$, it is crucial to accurately measure the linear velocity of the end-effector of the robot in real time. Without additional sensors on the end-effector, the linear velocity of the end-effector can be calculated by using the joint velocities in the joint space and the Jacobian matrix. However, in tendon-driven mechanisms like endoscopic robots, it is not possible to attach sensors such as encoders to each joint. Moreover, due to the hysteresis problem [49], the commanded motion at the motor may not be perfectly applied to each joint of the robot, which can lead to low accuracy in the calculated linear velocity of the end-effector. Therefore, by integrating a method capable of overcoming hysteresis with nonlinear characteristics, such as the Bouc-Wen model [50], for estimating the linear velocity of the end-effector, ideal results similar to those in the simulation environment are expected to be achieved in the actual operating environment of the robotic system.
As explained in Section III-A, a marking process should be performed before the cutting process for tumor removal. As indicated by (18) – (22), this study assumes that all information, such as marking points, tumor sizes, and locations, is provided beforehand. However, in actual clinical environments, the geometric information and locations of the tumors should be estimated from the images inputted from the endoscopic camera mounted on the robot system. To enable the proposed method to be used in real clinical environments, additional algorithms need to be integrated into the framework. For example, using stereo matching algorithms like RAFT-Stereo [51] and segmentation algorithms such as the MedSAM [52] would allow for three-dimensional analysis of objects in the task environment. This approach would enable the estimation of tumor shapes, sizes, and locations.
@@
Additionally, the interaction between the robot and tissue could happen and lead to external forces in real-world experiments. Although we did not consider interaction forces with tissues when cutting tumors in the simulations, the interaction forces could be dealt with as one of the disturbances $\bm{d}$ in the controller. Therefore, the controller used in the proposed method, (1), also ensures robust safety against external forces arising from these interactions.
@@ VII Conclusion and Future Work
In future work, the proposed framework will be improved to extend beyond simulations, enabling the robot system to be applied in real clinical environments. In particular, by applying a hysteresis model capable of representing non-linearities, the inherent issues of tendon-driven systems will be addressed, improving the accuracy of end-effector linear velocity estimation. Furthermore, through the integration of the endoscopic camera and deep learning algorithms, it will be possible to acquire geometric information of tumors in the task environment.
Acknowledgment
@@ Appendix A Author Information
[Uncaptioned image] Hansoul Kim received the B.S. degree from the Division of Robotics, Kwangwoon University, Seoul, Republic of Korea, in 2017, the M.S. degree from the Robotics Program, Korea Advanced Institute of Science and Technology (KAIST), Daejeon, Republic of Korea, in 2019, and the Ph.D. degree from the Department of Mechanical Engineering, KAIST, in 2022. From 2022 to 2023, he served as a Post-doctoral researcher at the Walker Department of Mechanical Engineering and Texas Robotics, the University of Texas at Austin, TX, USA. From 2023 to 2024, he served as as a Post-doctoral researcher in the Department of Electrical Engineering and Computer Sciences at the University of California, Berkeley, CA, USA. He is currently an Assistant Professor with the Department of Mechanical Engineering at the Myongji University, Yongin, Republic of Korea. His research interests primarily focus on surgical robotics, AI-assisted robot surgery, robot task automation, soft robotics, and mechanism design.

### v2

@@ VI Discussion
This study proposed a novel framework that can ensure the safety of automation for robotic surgery in a theoretically rigorous way. Our results were demonstrated in a scenario based on ESD, a representative endoscopic surgical method for the treatment of early-stage GI cancer, by using an endoscopic robot. The proposed framework utilizes CBFs to enable the automated motion of the robot, allowing it to clearly respect the boundaries of each tumor, even when multiple tumors are in close proximity within the GI tract. This allows individual removal of each tumor while enforcing safety guarantees on the surrounding normal mucosal tissues.
Although this study focuses on ensuring the safety of cutting task automation during ESD, extending the application to various tasks such as suturing or clamping by including not only the definition of the safe set boundary but also a model that accounts for the nonlinear interactions between superelastic tissues and the robot is required to bring the framework closer to practical clinical application. Additionally, considering superelastic characteristics, such as the texture and stiffness of the tumor, are considered, is needed to validate the robustness of the framework to variability that may arise in real-world applications.
@@
A model-free control scheme was leveraged, thus avoiding the complexity of robot dynamics modeling. However, even though a model-free control scheme was used, the simulation environment can differ from the actual environment where the robotic system is implemented. From (24), the proposed framework employs a simple model-free velocity controller to track the safe velocity $\dot{\bm{x}}_{s}$. Here, since $\dot{\bm{e}}$ is expressed as $J^{+}(\bm{q})(\dot{\bm{x}}-\dot{\bm{x}}_{s})$, it is crucial to accurately measure the linear velocity of the end-effector of the robot in real time. Without additional sensors on the end-effector, the linear velocity of the end-effector can be calculated by using the joint velocities in the joint space and the Jacobian matrix. However, in tendon-driven mechanisms like endoscopic robots, it is not possible to attach sensors such as encoders to each joint. Moreover, due to the hysteresis problem [49], the commanded motion at the motor may not be perfectly applied to each joint of the robot, which can lead to low accuracy in the calculated linear velocity of the end-effector. Therefore, by integrating a method capable of overcoming hysteresis with nonlinear characteristics, such as the Bouc-Wen model [50], for estimating the linear velocity of the end-effector, ideal results similar to those in the simulation environment are expected to be achieved in the actual operating environment of the robotic system. As an alternative to hysteresis modeling, a system with two motors per bending degree of freedom that mechanically compensates for hysteresis by minimizing slack could be adopted [51].
As explained in Section III-A, a marking process should be performed before the cutting process for tumor removal. As indicated by (18) – (22), this study assumes that all information, such as marking points, tumor sizes, and locations, is provided beforehand. However, in actual clinical environments, the geometric information and locations of the tumors should be estimated from the images inputted from the endoscopic camera mounted on the robot system, or registered during the pre-operative examination stage. To enable the proposed method to be used in real clinical environments, additional algorithms need to be integrated into the framework. For example, using stereo matching algorithms like RAFT-Stereo [52] and segmentation algorithms such as the MedSAM [53] would allow for three-dimensional analysis of objects in the task environment. This approach would enable the estimation of tumor shapes, sizes, and locations. The method proposed in this study focuses on ensuring the safety of automated robotic motion. When combined with the aforementioned deep learning-based vision algorithms for precise tumor localization, the proposed framework is expected to contribute to the ongoing advancement of medical robotics.
@@
Additionally, unexpected external forces in the real-world experiment could happen due to the uncertainty of the system model such as the robot model, deformability of the body tissues (e.g. GI), and the interaction forces between the robot and tissues, which could lead to safety violations. However, the unexpected external forces could be dealt with as one of the disturbances $\bm{d}$ in the controller as represented in (16). Although we did not consider those uncertainties when cutting tumors in the simulations conducted in this paper, it is possible to achieve robust safety against unexpected external forces arising from uncertainties, since we use the robust CBF-based controller with input-to-state stability and safety conditions, based on Corollary 1. Alternatively, we can ensure safety guarantees while reducing unknown uncertainty simultaneously by estimating parametric and non-parametric uncertainty based on adaptive CBF[55], robust adaptive CBF [36], Gaussian Process CBF[56], Gaussiance Process robust adaptive CBF[38], disturbance observer-based CBF [57][58], and several neural network-based CBFs [59] [60] [61].
Lastly, tuning the $\alpha$ value is non-trivial, as the parameter depends on model uncertainty and the size of the safe region, while significantly affecting the performances of CBF-based controllers. For example, when the safe region is very small and there exists uncertainty in the system model, CBF-based controllers might not be able to find the feasible safe control input unless $\alpha$ is not properly selected. To address this issue, we can use the sampling-based approach for tuning $\alpha$ proposed by [62]. First, we sample states that satisfy the hard safe constraint condition, for example, $\dot{h}\geq\alpha(h)$. Using the sampled states, we can construct the approximated safe set and consequently verify whether the selected $\alpha$ is valid. We expect that an iterative tuning algorithm for $\alpha$ can be incorporated to enhance the applicability of the proposed framework in real-world experiments in future research.
@@ VII Conclusion and Future Work
In future work, the proposed framework will be further developed to move beyond simulations, enabling its application to realistic phantom models or ex-vivo environments. In particular, the focus will be on developing a method to estimate the linear velocity of the end-effector in a tendon-driven robot based on motor rotational speed, utilizing the robot’s kinematics and a nonlinear hysteresis model for real-world applications of the proposed control framework. Furthermore, through the integration of the endoscopic camera and deep learning algorithms, it will be possible to acquire geometric information of tumors in the task environment.
@@ Appendix A Author Information
[Uncaptioned image] Hansoul Kim received the B.S. degree from the Division of Robotics, Kwangwoon University, Seoul, Republic of Korea, in 2017, the M.S. degree from the Robotics Program, Korea Advanced Institute of Science and Technology (KAIST), Daejeon, Republic of Korea, in 2019, and the Ph.D. degree from the Department of Mechanical Engineering, KAIST, in 2022. From 2022 to 2023, he served as a Post-doctoral researcher at the Walker Department of Mechanical Engineering and Texas Robotics, the University of Texas at Austin, TX, USA. From 2023 to 2024, he served as as a Post-doctoral researcher in the Department of Electrical Engineering and Computer Sciences at the University of California, Berkeley, CA, USA. He is currently an Assistant Professor with the Division of Mechanical System Engineering at the Myongji University, Yongin, Republic of Korea. His research interests primarily focus on surgical robotics, AI-assisted robot surgery, robot task automation, soft robotics, and mechanism design.